### v1

# REST API Testing in DevOps: A Study on an Evolving Healthcare IoT Application

Abstract
Healthcare Internet of Things (IoT) applications often integrate various third-party healthcare applications and medical devices through REST APIs, resulting in complex and interdependent networks of REST APIs. Oslo City’s healthcare department collaborates with various industry partners to develop such healthcare IoT applications enriched with a diverse set of REST APIs. Following the DevOps process, these REST APIs continuously evolve to accommodate evolving needs such as new features, services, and devices. Oslo City’s primary goal is to utilize automated solutions for continuous testing of these REST APIs at each evolution stage, thereby ensuring their dependability. Although the literature offers various automated REST API testing tools, their effectiveness in regression testing of the evolving REST APIs of healthcare IoT applications within a DevOps context remains undetermined. This paper evaluates state-of-the-art and well-established REST API testing tools—specifically, RESTest, EvoMaster, Schemathesis, RESTler, and RestTestGen—for the regression testing of a real-world healthcare IoT application, considering failures, faults, coverage, regressions, and cost. We conducted experiments using all accessible REST APIs (17 APIs with 120 endpoints), and 14 releases evolved during DevOps. Overall, all tools generated tests leading to several failures, 18 potential faults, up to 84% coverage, 23 regressions, and over 80% cost overhead.
Keywords Black-box Testing, Fuzzing, REST API, Web Services, DevOps, CI/CD, Healthcare IoT
1 Introduction
Healthcare Internet of Things (IoT) applications commonly use a cloud-based architecture as a hub for stakeholders such as healthcare professionals, caretakers, and patients [1]. This involves integrating various healthcare applications serving different purposes (such as electronic health records (EHR) for managing patient data) and medical devices (like medicine dispensers) assigned to patients. These integrations into the IoT cloud are facilitated through Application Programming Interfaces (APIs) following the Representational State Transfer (REST) architecture—concisely REST APIs. A malfunction in REST APIs can significantly dismantle the healthcare IoT cloud, potentially making crucial healthcare services unavailable to patients. This could fail to provide timely healthcare services, which can have serious consequences. Given the critical nature of healthcare IoT applications, rigorous testing at various levels, such as the system level, is essential to ensure their dependability.
Oslo City’s healthcare department [2] collaborates with various healthcare industries to develop IoT applications, aiming to deliver efficient healthcare services to residents. These collaborations typically adhere to DevOps practices, where the Oslo City healthcare team continuously manages, analyzes, and reports problems within the IoT healthcare application under development. One of their activities is to test the REST APIs of such an application at the system level with integrated medical devices and third-party applications. However, manual testing of REST APIs at each evolution stage can be time-consuming and laborious due to the complex and interdependent nature of these APIs, thus necessitating an automated testing solution.
Over time, automated testing of REST APIs has attracted significant interest. Numerous approaches, such as RESTest [3], RESTler [4], Schemathesis [5], RESTCT [6], RestTestGen [7], and EvoMaster [8], have been proposed. The effectiveness of REST API testing approaches has been examined in different empirical studies (e.g., [9, 10, 11]) for various open-source and industrial web applications. Our analysis of the literature revealed two key observations. Firstly, the effectiveness of REST API testing approaches in testing IoT applications’ REST APIs remains unexplored. Secondly, the use of REST API testing approaches in regression testing of evolving REST APIs within the DevOps context is also unexplored. We found only one study by Godefroid et al. [12] that employed RESTler for regression testing of open-source APIs using historical release data. Thus, the evaluation of REST API testing approaches in regression testing of continuously evolving REST APIs of real-world IoT applications remains a research gap. Considering this gap, we assess well-established approaches for testing the REST API of a real-world healthcare IoT application deployed in Oslo City and analyze their practicality in this context.
This paper builds upon our previous work [13] in which we reported an empirical evaluation of RESTest for testing six APIs with 41 endpoints of a healthcare IoT application’s production release. This paper presents a large-scale empirical study that evaluates state-of-the-art and well-established tools—specifically RESTest, EvoMaster, Schemathesis, RESTler, and RestTestGen—for the regression testing of evolving REST APIs associated with a real-world healthcare IoT application, which is progressively developed via the DevOps process. Our experiments encompass 14 different types of releases (daily, weekly, and monthly) and feature all accessible REST APIs of the application under test, i.e., 17 APIs with 120 endpoints. We evaluated REST API testing tools, considering their effectiveness in terms of failures, faults, coverage, regressions, and cost overhead. Results show that EvoMaster generated the highest number of client-side and server-side failures among all the tools. The fault analysis revealed that the tests generated by all tools led to various potential faults, totaling 18 unique faults across 14 releases. The coverage results indicate that RESTest achieved approximately 82% coverage, while EvoMaster, Schemathesis, RESTler, and RestTestGen attained an overall coverage of around 84%. Moreover, analysis of regressions across all 14 releases demonstrated that the tests generated by each tool resulted in the identification of several potential regressions, with EvoMaster at the top. In total, 23 unique regressions were discovered through testing results from all tools across the 14 releases. Through these regressions, we analyzed the stability of APIs during continuous releases, specifically focusing on the state of API implementation (i.e., feature upgrades) and the propagation of failures. The results of this regression analysis offer valuable insights for practitioners, assisting them in identifying potential areas of API vulnerability, determining recurring failures, and choosing a suitable testing tool. Finally, the cost analysis demonstrated that each tool generates a significant number of tests, often leading to numerous duplicate failures, thereby resulting in a cost overhead exceeding 80%.
The specific contributions of this work are summarized below:
  • A large-scale empirical study for testing REST APIs of a real-world healthcare IoT application featuring:
      Five REST API testing tools, i.e., RESTest, EvoMaster, Schemathesis, RESTler, and RestTestGen configured with different black-box testing techniques, making a total to 10 tools.
      A total of 17 REST APIs comprising 120 endpoints.
      14 rapid releases (including daily, weekly, and monthly) of our industrial healthcare IoT application continuously built in a staging environment following DevOps practices.
  • A guide for practitioners on selecting REST API testing tools, derived from our empirical findings.
  • Experiences, insights, and challenges for researchers and practitioners.
The remainder of the paper is organized as follows. An overview of healthcare IoT applications and our industrial context are described in Section 2. Related works are discussed in Section 3. The experiments are reported in Section 4. Recommendations for selecting REST API testing tools are presented in Section 5. Experiences, insights, and challenges are outlined in Section 6. The paper concludes in Section 7.
2 Background
We provide an overview of healthcare IoT applications, our industrial context, and the associated challenges.
2.1 Healthcare IoT Applications
A typical layout of a healthcare IoT application is depicted in Figure 1. A healthcare IoT application utilizes a cloud platform that offers web and mobile interfaces for various users, including medical professionals, healthcare authorities, caregivers, and patients. Within the IoT cloud, it links to numerous third-party applications that provide necessary healthcare services, such as EHR and pharmacies. It is also connected to various medical devices assigned to patients based on specific healthcare needs, such as medicine dispensers for timely medication and pulse oximeters for heart rate monitoring.
All communications with third-party applications and medical devices are facilitated through REST APIs. Each service provider, whether for a device or an external application, has its own server and API implementation. The REST APIs of a healthcare IoT application communicate with the APIs of medical devices and external applications using the JavaScript Object Notation (JSON) format for data exchange. Consequently, the healthcare IoT application possesses a rich set of REST APIs with complicated structures due to their dependencies on the APIs of multiple types of applications. For example, in a patient medication scenario, the healthcare IoT application initially retrieves a medication plan via an API request to the pharmacy’s API. It then interprets this medication plan and updates the web interface for the healthcare personnel. Finally, it sends the medication plan to the patient’s medicine dispenser via another API call to the device’s API.
Figure 1: A typical layout of a healthcare IoT application and the focus of our work
2.2 Industrial Context and Challenges
The Directorate of Health in Norway launched the national welfare technology program [14] to offer efficient digital home monitoring and healthcare services. As part of this program, Oslo City’s healthcare department [2] worked with different industry partners to create a cloud-based platform with an integrated network of smart medical devices and healthcare service providers — a platform that resembles the layout of the IoT healthcare applications described in Section 2.1. Oslo City follows the DevOps process to foster these collaborations with industry partners. In this process, Oslo City’s healthcare team continuously manages, analyzes, and reports any issues that arise within the under-development healthcare IoT application. Figure 2 illustrates the commonly followed DevOps process for healthcare IoT applications. In this process, Oslo City initially gets involved during the Plan phase, where they define system requirements and plan various project aspects. The subsequent phases, including Code, Build, Test, Release, and Deploy, involve contributions from industry partners. Upon deploying a specific release, Oslo City’s healthcare team handles the Operate and Monitor phases, in which they analyze the delivered functionalities to ensure dependability and report any issues or necessary enhancements.
Figure 2: An overview of the DevOps process for the healthcare IoT application in our context, highlighting the phases involving industry partners, the involvement of Oslo City, and the specific focus of our paper.
Throughout the DevOps process, healthcare IoT applications continuously evolve, adding new services or upgrading existing ones. REST APIs, which make up a significant part of these applications’ backend, also undergo regular evolution. These stages of evolution typically correspond to daily builds for minor changes or bug fixes and weekly builds for major updates or new features. Testing REST APIs at each stage of evolution is labor-intensive and time-consuming. Therefore, automated testing of REST APIs at each evolution stage is essential to ensure the dependability of healthcare IoT applications.
Automated REST API testing of evolving REST APIs presents several challenges. The requirement for frequent testing at each stage of evolution, especially when connected to medical devices and third-party applications, may risk damaging the device or causing service blocking from external applications [1, 15]. Additionally, identifying the root causes of faults or failures due to the dependencies of REST APIs on various third-party applications is a considerable challenge [13]. For testing REST APIs of healthcare IoT applications, generating test data requires specific knowledge of the healthcare domain and an understanding of the application context. Lastly, the complexity and interdependent structure of REST APIs make automated test oracle generation a challenging task. These challenges were primarily addressed in our previous works [15, 16, 17]. In this study, we aim to evaluate REST API testing tools within the context of an evolving healthcare IoT application and to determine the most effective tools for integration into the overall test infrastructure (as indicated in Figure 1).
3 Related Works
We present related works on testing IoT applications, REST API testing, and empirical evaluations of REST API testing tools.
3.1 IoT Applications Testing
Numerous studies focus on various aspects of testing IoT applications [18]. These include model-based conformance testing of the IoT [19], user interface-driven acceptance testing of IoT systems [20], testing the heterogeneity of IoT devices in the loop [21], and identifying faults in the integration of devices with IoT applications [22]. Other research areas include combinatorial testing and coverage criteria for IoT systems [23], combinatorial testing for IoT-based smart home applications [24], and simulating the health monitoring activities of healthcare IoT applications [25]. In contrast to the studies mentioned above, our work primarily focuses on the regression testing of evolving REST APIs within a healthcare IoT application. Furthermore, our research involves the empirical evaluation of REST API testing tools, as opposed to proposing a new testing technique. The insights derived from our empirical study could potentially be employed to enhance existing techniques for testing IoT applications across various domains.
3.2 REST API Testing
Automated testing of REST APIs has gained significant interest, and many approaches have been proposed to address this [26]. These comprise RESTest [3, 27], ARTE [28]—an extension to RESTest, RESTler [4, 29], Schemathesis [5], RESTCT [6], RestTestGen [7, 30], EvoMaster [8, 31], RapiTest [32], QuickREST [33], NLP2REST [34], and ARAT-RL [35]. Some open-source tools like APIFuzzer [36], Tcases [37], and Dredd [38] also support testing of REST APIs. However, as reported in [9], these tools did not outperform research-based tools. In addition, numerous studies target various forms of REST API testing, including model-based testing [39], specification-based testing [40], robustness testing [41], metamorphic testing [42], search-based test case improvement [43], security testing [44], test input validation using deep learning [45], and identifying vulnerabilities in REST APIs [46]. Although these studies propose various approaches for testing REST APIs, our work explicitly evaluates selected tools for testing the REST APIs of a real-world healthcare IoT application within a DevOps environment. In another relevant work, Godefroid et al. [12] proposed an approach for regression testing of REST APIs and evaluated their approach using the RESTler tool on open-source APIs. Compared to this work, our work differs in two aspects. First, our analysis focuses on assessing REST API testing tools in a continuous regression testing environment, whereas their analysis is based on historical release data. Second, we utilized five REST API testing tools along with their supported testing techniques (amounting to a total of 10 tools), in contrast to their work which only considered one tool, specifically RESTler.
3.3 Empirical Evaluations of REST API Testing Tools
Several studies have conducted empirical research comparing various REST API testing tools. In this regard, Martin et al. [10] compared black-box and white-box testing tools, specifically RESTest and EvoMaster, by using them in four open-source APIs. Corradini et al. [47] also delved into this subject, analyzing four tools: RESTest, RESTler, RestTestGen, and bBOXRT. Their study involved testing the REST APIs of 14 open-source applications. Furthermore, Kim et al. [9] evaluated 10 different tools designed for testing REST APIs, using 20 open-source applications in their study. More recently, Zhang et al. [48] performed an empirical comparison of seven REST API testing tools using 20 different case studies, which included an industrial case study. While the studies outlined above assess various REST API testing tools across different case studies, two key aspects remain unexplored. The first is the effectiveness of these testing tools when applied to REST APIs of large-scale IoT applications. The second is the applicability of these tools for regression testing of evolving REST APIs within a DevOps context. Our work specifically targets these aspects.
In our previous work [13], we evaluated the effectiveness of RESTest in testing six REST APIs of an IoT application in industrial healthcare. In this paper, we extend this scope by evaluating five REST API testing tools for testing continuously evolving REST APIs of a real-world healthcare IoT application in a DevOps context. The specific additions to previous work are outlined below.
  • We incorporate an additional 11 REST APIs with 79 endpoints of a healthcare IoT application targeting different features, resulting in a total of 17 APIs with 120 endpoints.
  • In addition to RESTest, we include four other REST API testing tools in our study, namely EvoMaster, Schemathesis, RESTler, and RestTestGen. Each tool is configured with its supported testing technique, resulting in 10 tools utilized in our experiment.
  • We evaluate REST API testing tools for continuous regression testing by utilizing 14 API releases (daily, weekly, and monthly) progressively built in a staging environment following DevOps practices.
  • We extended the evaluation scope by introducing two additional research questions: one focusing on regression analysis and the other on cost overhead analysis.
  • We present recommendations to assist practitioners in selecting appropriate tools. Furthermore, based on our large-scale empirical evaluation, we provide valuable insights and lessons for researchers and practitioners.
4 Experiments
This section presents experiments to evaluate REST API testing tools in a real-world healthcare IoT application context. Our evaluation primarily focuses on regression testing of REST APIs that progressively evolved during the DevOps process. Specifically, we analyze REST API testing tools regarding failures, faults, coverage, regressions, and cost overhead. Given this focus, we formulate the following research questions (RQs).
  • RQ1:
    How does the failure detection ability of REST API testing tools vary as APIs under test evolve?
    As APIs undergo upgrades with new/enhanced features, it becomes essential to ensure that the REST API testing tools can detect any API failures arising from these changes. The findings from this RQ can assist practitioners in selecting appropriate tools for testing in the continuous integration and deployment phases of the DevOps process.
  • RQ2:
    What kinds of potential faults can be discovered using REST API testing tools during API evolution?
    Faults in APIs, which can result in API failures, are important to investigate, especially in critical systems like healthcare. The potential faults identified in this RQ could assist practitioners in fixing API implementations. Furthermore, they can guide practitioners during API development to avoid frequently occurring faults throughout the API’s evolution.
  • RQ3:
    What level of API coverage does each REST API testing tool achieve across various API releases?
    API coverage serves as an indicator of the effectiveness of the test cases generated by a testing tool. Therefore, the coverage analysis in this RQ demonstrates the API coverage that each REST API testing tool achieves during test execution. These findings could assist practitioners in selecting tools that consistently exhibit high and stable coverage across multiple evolving releases.
  • RQ4:
    What types of potential regressions can be uncovered from the outcomes of various REST API testing tools?
    In this RQ, we investigate the test results generated by each REST API testing tool to find potential regressions in evolving API releases.
  • RQ5:
    What are the cost implications of using different REST API testing tools?
    Testing frequent and rapid releases in DevOps could involve substantial costs, such as overheads associated with API calls. From this RQ, we evaluate the relative cost-effectiveness of various REST API testing tools. This can guide decision-making regarding which tool to adopt, based on a balance of cost and testing effectiveness.
4.1 REST APIs Under Test
Table 1 presents the characteristics of all accessible APIs [^1] APIs for which the access provided by Oslo City as a part of experimental apparatus for research purposes. Inaccessible APIs are for sensitive data with privacy concerns. used in our experiments. Each API is associated with a specific functionality within the healthcare IoT application under test. These APIs consist of multiple endpoints, each targeting different types of features. The Alerts API has five features for assigning alerts, closing alerts, all alerts, total number of alerts, and dashboard alerts. The Authentication API targets different methods for authenticating different users. This includes authentication through email, password, SMS, phone number, device’s IMEI, token, medical device, and login credentials. In addition, it includes features such as password recovery for forgotten passwords, updating login credentials, and creating an emergency profile for a user. Three device APIs are related to three medicine dispensers (i.e., Karie [49], Medido [50], and Pilly [51]) that are integrated with the healthcare IoT application. Karie API provides features for modifying device settings, creating medication plans, and loading medical plans from pharmacies. Medido API enables modifying device settings and creating mock and real medication plans. Pilly API only allows for the configuration of device settings.
The API for Patients targets features are related to patients’ health information, billings, time logs, health tracking, patients’ actions to perform, and reimbursements. The Users API focuses on user actions, that is, adding notes, creating events, threshold, searching users based on name or birth date, validating a user, updating user information, counting or listing down users, logging user interactions, and matching with existing patients. The Courses API is concerned with e-learning courses with features for course category, courses list, course modules, linking a course, uploading a new course, and finishing a course. The API for Mobile App targets features related to app version, app information, patient’s timeline, password information, and checking old and new passwords. The Reimbursements API includes features for creating billing summaries, reimbursements time logs, patient reimbursements status, measurements related reimbursements, task identifiers, and tracking reimbursements.
The API for User Tasks provides features for various user tasks, i.e. assigning tasks to the concerned users, patient health surveys, task lists, task reminders, general tasks for administrative purposes, and global surveys for different users. The Invoicing API provides features for adding, canceling, replacing, terminating, and subscribing invoices. The Catalog API supplies catalogs with various types, models, units, features, and lists of catalogs. The API for Reports provides reports in different formats, patients’ reports, patients’ medication reports, zone-wise reports, and event-specific reports. The Care API is related to patient care messages, events, and reminder status. The Development API is concerned with creating smart alert rules, user profiles, and temporary interface generation.
Table 1: Characteristics of APIs selected for the experiments, i.e., API label, number of endpoints (EPs), and features
API Label EPs Features
Alerts 5 Assign, Close, All, Total, Dashboard
Authentication 13 Email, Password, SMS, Phone, IMEI, Token, Device, Log-in, Forgot password (Send/Resend), Credentials, Emergency profile
Device-Karie 3 Settings, Plan, Pharmacy
Device-Medido 3 Settings, Plan, Mock Plan
Device-Pilly 1 Settings
Patients 6 Health data, Billings, Time logs, Tracking, Actions, Reimbursements
Measurements 5 Single & Multiple inputs (Manual/Auto), Time series
Users 14 Notes, Events, Threshold, Search methods, Validate, Update, Count/List, Logs, Match patient
Courses 6 Category, List, Module, Link, Upload, Finish
Mobile App 6 Version, Information, Timeline, Password info, Check old & new password
Reimbursements 11 Billing summary, Time logs, Patients, Measurements, Task identifiers, Track
User Tasks 9 Assignments, Survey, Tasks, Reminder, General tasks, Global survey
Invoicing 9 Add, Cancel, Replace, Terminate, Subscriptions (Check, Fix, Aggregate)
Catalog 8 Types, Models, Units, Features, List
Reports 10 Formats (HTML, JSON, PDF), Patients, Medication, Zone, Events
Care 4 Messages, Events, Reminder status
Development 7 Smart alert rules, Profiles, Auto UI generation
Total (APIs: 17) 120 -
4.2 Testing Tools Selection
Table 2 presents information for REST API testing tools selected for our experiments, including details such as tool name, version, release date, and applied testing techniques. The most recent version of each tool was picked as of May 10, 2024. To select tools for our experiments, we first analyzed empirical studies focusing on tool comparison. Drawing from the recommendations of effective and widely-used tools in previous studies [26, 48, 9], we initially shortlisted eight tools for our experiments, including RESTest, EvoMaster, RESTler, RESTCT, RestTestGen, bBOXRT, Morest, and Schemathesis. We later expanded our list to include TnT-Fuzzer, Got-Swag, RapiTest, QuickREST, and three recent tools: NLPtoREST, ARAT-RL, and foREST. Initially, we configured all tools to test the REST APIs of our application under test. Upon running these tools, we encountered various compatibility issues and crashes. After diagnosing these problems, we identified their common causes. Based on our investigation and insights from previous empirical studies [48, 9], we formulated a set of criteria required for a tool to qualify for testing our application under test. Note that resolving issues related to the tools could potentially introduce an internal validity threat, which also falls outside the scope of our paper. Next, we outline the criteria and elaborate on our systematic approach to the tool selection process.
  • C1.
    Tools that support OpenAPI specification version 3.0, ensuring compatibility with the REST APIs of our healthcare IoT application under test.
  • C2.
    Well-established tools that have undergone continuous development phases, showcasing a commitment to improvement through the addition of new features, bug fixes, and the release of stable versions.
  • C3.
    Tools applied to numerous cross-domain industrial systems and support execution on multiple platforms.
  • C4.
    Research-based tools derived from various research works with comprehensive and well-founded techniques (similar to earlier works [48, 9]).
Among all available REST API testing tools, RESTCT only supports the OpenAPI specification version 2.0 (violating C1). Tools like TnT-Fuzzer and Got-Swag, as referenced in [5], also support OpenAPI version 2.0, thereby violating C1. The public version of Morest only provides a replication package, and its repository was last updated on September 13, 2021 (violating C2). bBOXRT provides an initial version as a replication package, last updated on January 30, 2020 (violating C2). In addition, bBOXRT requires an input API file in Java, which is not available in our case. In a similar line of work, RapiTest and QuickREST only provide replication packages, violating C2. Recent tools NLPtoREST and ARAT-RL provide replication packages, in the initial development stages, and can be run on Ubuntu (violating C2 & C3). The tool link for another recent black-box testing tool foREST [52] is also unavailable. Moreover, foREST is in early development phases, violating C2. Tools like Tcases, APIFuzzer, and Dredd are open-source tools without a research background, violating criterion C4. After careful inspection, we selected RESTest, EvoMaster, Schemathesis, RESTler, and RestTestGen for our experiments. Below is a brief overview of these selected REST API testing tools.
RESTest [3] is an automated tool designed to test REST APIs of web applications, which work mainly in a black-box testing mode. This tool utilizes the API specification and test configurations as input to generate test cases. For test generation, RESTest supports various testing techniques, including constraint-based testing (CBT), adaptive random testing (ART), random testing (RT), and fuzz testing (FT). In addition, RESTest offers advanced features such as online testing in a production environment, handling inter-parameter dependencies, and generating realistic test data.
EvoMaster [8] is a Web API testing tool that employs evolutionary techniques to generate test cases. It supports black-box and white-box testing modes and is compatible with various API types, including REST and GraphQL. In black-box testing, EvoMaster takes the API specification as input. White-box testing requires additional inputs, including source code and a driver. For black-box mode, EvoMaster employs fuzzing techniques for test generation, while in white-box mode, it utilizes many objective search algorithms for test generation.
Schemathesis [5] is a web API fuzz testing tool that utilizes property-based testing for generating test cases. It operates in black-box and white-box testing modes and is compatible with APIs specified in OpenAPI or GraphQL formats. This tool is designed to test cross-platform web APIs, irrespective of the programming language used for development. In addition, Schemathesis has been used in various industrial projects, thereby streamlining the process of automated REST API testing.
RESTler [4] is a web services fuzz testing tool mainly focusing on identifying security and reliability faults in REST APIs. It takes API specification in OpenAPI format, analyzes the specification to identify dependencies among requests, generates fuzz tests, and executes tests on the API to explore web services. Based on responses originating in various test executions, it learns distinct bug patterns. It supports two types of fuzzing, i.e., fuzz and fuzz-lean.
RestTestGen [7] is a REST API testing tool with a primary focus on black-box testing. It supports two testing techniques: nominal and error testing and mass assignment security testing. RestTestGen offers an extensible interface that facilitates adding new features and implementing novel testing techniques. In this context, a recent tool NLP2REST [34] is integrated with RestTestGen, leveraging natural language processing techniques to enhance functionality in API testing scenarios.
Table 2: REST API testing tools used in the experiments, comprising details like the tool’s name, version, release date, and applied testing techniques
Tool Version Release Date Testing Techniques
RESTest 1.4.0 Nov. 12, 2023 CBT, ART, RT, FT
EvoMaster 3.0.0 Apr. 05, 2024 Fuzzing
Schemathesis 3.27.1 Apr. 29, 2024 Property-based testing
RESTler 9.2.4 Mar. 06, 2024 Fuzz, Fuzz-lean
RestTestGen 24.09 Mar. 30, 2024 Nominal and error testing, Mass assignment security testing
Figure 3: Workflow of experiments design, setup, and execution for testing multiple releases of REST APIs, a healthcare IoT application continuously developed during the DevOps process.
4.3 Experiments Design
We designed our experiments to test the REST APIs of a real-world healthcare IoT application that continually evolved in DevOps. In the following, we elaborate on the details of the preparation and execution of the experiment.
4.3.1 Experiment Preparation
Our experiments involve five REST API testing tools, each supporting one or more testing techniques. For each testing mode, we set up different configurations of these tools. RESTest configured with the CBT, ART, RT and FT testing techniques is termed RESTest-CBT, RESTest-ART, RESTest-RT and RESTest-FT, abbreviated as RT-CBT, RT-ART, RT-RT and RT-FT, respectively. EvoMaster configured with back-box testing mode is termed EvoMaster-BB (in short EM-BB). Schemathesis configured with property-based testing mode is termed Schemathesis-PT (in short ST-PT). RESTler configured with fuzz and fuzz-lean testing modes is termed RESTler-F and RESTler-FL, abbreviated as RL-Fz and RL-FL, respectively. RestTestGen configured with nominal and error testing and mass assignment security testing is termed RestTestGen-NET and RestTestGen-MAST, abbreviated as RTG-NET and RTG-MAST respectively. These configurations result in a total of 10 tools for our experiments.
We set up our experiments based on information from Oslo City’s industry partner about planned releases, i.e., daily, weekly, and monthly. Our experiments were set up to use 14 distinct releases of our application under test, denoted as R1 through R14. This includes eight daily and five weekly builds in the staging environment and one monthly production release. The releases were structured as follows: R1, R4, R7, R10, and R13 were weekly releases; R2, R3, R5, R6, R8, R9, R11, and R12 were daily releases; finally, R14 was a monthly release. Note that the specific dates corresponding to the releases on which we conducted our experiments are not disclosed to maintain the confidentiality of Oslo City’s industry partner.
Figure 3 shows the design of the experiment. We started with the first weekly release (R1) and ran all testing tools on that release. At the end of execution, we collected test cases and results produced by each testing tool for R1. For each next release, R2-R14, we executed the same set of tests generated by each tool on each subsequent release. One of the primary objectives of our experiment is to identify regressions. To achieve this, using the same set of tests (generated from each tool) is necessary. This is due to the inherent randomness factor involved in each testing tool, leading to potential variability in the tests generated at different times. For example, the tests generated during the initial run may differ from those generated during subsequent runs. This variability could hinder the detection of regressions in evolving releases. Therefore, to effectively identify regressions across different releases, we used the set of tests generated for R1 to test releases R2-R14. This approach to regression testing is commonly referred to as the retest-all strategy in literature [53, 54]. We chose this strategy because it does not necessitate a test selection, minimization, or prioritization approach to determine which specific test cases should be rerun. To the best of our knowledge, no such approach exists for REST API test cases. Furthermore, adapting existing test selection, minimization, or prioritization approaches for test cases created for REST APIs could potentially introduce significant bias, such as implementation bias [16]. This also deviates from the primary focus of our study.
4.3.2 Experiment Execution
We scheduled our experiments for each daily build to run overnight. For the weekly release, experiments were set to run over the weekend. The experiment for the monthly release was also arranged to run overnight, considering they were scheduled during the workweek. The overall duration of our experiments was more than a month, specifically lasting 4 weeks and 5 days. We executed all experiments on one machine with specifications of a four-core 3.6 GHz CPU, 32 GB RAM, and Windows 10 operating system.
To account for the inherent randomness of REST API testing tools, we executed each tool 10 times in parallel, with each run limited to one hour. Furthermore, our experiments involve three real medical devices in the loop. Given the constraints on the ability to handle requests from these devices and the potential damage to the device [1, 15], we introduced a delay of three seconds for APIs that require interaction with medical devices. This execution setup aligns with that used in our previous work [13].
Throughout the execution of our experiments, we continuously compiled tests generated by each REST API testing tool. This data comprised comprehensive details of the requests generated by each testing tool and the corresponding responses returned from the APIs under test. At the end of each experiment run, we collected and reviewed the data, and then stored it in an organized manner corresponding to each release, tool, and API. The data collection process was designed to facilitate data analysis and interpretation for each RQ.
Table 3: Evaluation aspect and analyses metrics, measures, and statistical tests corresponding to each RQ
RQ Aspect Analysis Metrics, Measures, and Statistical Tests
RQ1 Failure Gestalt similarity, Unique failures, Wilcoxon test, Vargha-Delaney Â12 measure
RQ2 Fault Gestalt similarity, Unique faults
RQ3 Coverage Percentage coverage, Wilcoxon test, Vargha-Delaney Â12 measure
RQ4 Regression Gestalt similarity, Unique regressions
RQ5 Cost Percentage cost
4.4 Data Analysis and Interpretation
Table 3 presents analysis metrics, measures, and statistical tests for each RQ, along with its corresponding evaluation aspect. To analyze RQ1 results related to failures, we computed unique failures representing server-side (5XX) and client-side (4XX) errors, accounting for the possibility of multiple occurrences of the same error type. For this purpose, we first grouped the 5XX and 4XX errors separately. We then calculated the similarity between failure responses using the Gestalt approach for string matching proposed by Ratcliff and Metzener [55], implemented in a Python library named difflib [56]. Given that the responses obtained in our study contain data in JSON and HTML formats, we opted for the Gestalt string matching algorithm, considering its application for code comparison [57]. The Gestalt string matching algorithm calculates the similarity between two strings on a range from 0 to 1, where 0 indicates no similarity and 1 denotes a 100% similarity [57, 56]. A threshold of 0.5 (or 50%) serves as a balanced midpoint on this scale. To identify unique failure responses in our study, we set a similarity threshold of 0.5. Responses with a similarity score below 0.5 are considered to exhibit low similarity, while those with a score above this threshold are deemed to have high similarity.
To analyze RQ2 results concerning faults, we followed a semi-automatic approach inspired by Martin-Lopez et al. [27] to discover potential faults. Initially, we grouped unique failure and success responses based on their similarity score, calculated using the Gestalt algorithm with the same threshold set for RQ1. We chose to include the success responses in this process, as our previous work has shown that even successful responses can potentially lead to faults [58]. Next, we created clusters for different groups by parsing the JSON and HTML responses to obtain information about the failures, such as error messages. Lastly, we carefully inspected the resulting list of potential faults and assigned each a unique identifier (ID) to facilitate precise interpretation and reporting of these faults.
To analyze coverage results for RQ3, we used Restats [59] to calculate the coverage of test cases generated by each REST API testing tool. The coverage reported by Restats encompasses different aspects, including paths, operations, status classes, statuses, response types, request types, parameters, and parameter values. For our analysis, we evaluated the overall coverage achieved by each testing tool for all releases.
To analyze RQ4 results related to regressions, we first assembled responses into pairs of two consecutive releases, such as R1-R2, R2-R3, etc. Subsequently, we compared the response status codes of each pair (for example, R1 and R2) to analyze any discrepancy between the status codes of two consecutive releases. If two status codes differed, we further compared the detailed responses of each pair by calculating their similarity using the Gestalt algorithm with the same threshold set for RQ1. Finally, we thoroughly inspected the set of potential regressions and marked each with a unique identifier (ID) to differentiate regressions while reporting results.
In the cost analysis for RQ5, we initially computed the failure and fault detection rates and subsequently calculated the percentages of cost overhead. We utilized Equation 1 to calculate the failure detection rates, and Equation 2 for computing fault detection rates. In these equations, $N_{failures}$ represents the total number of unique failures, $T$ denotes the total number of tests generated by a tool, and $M_{faults}$ symbolizes the total number of unique faults identified from failure results. To compute the cost overhead for the rates corresponding to failures or faults, we use Equation 3. In this equation, we subtract the rates associated with failures or faults from 100. The primary objective is to determine the percentages of failed tests to identify unique failures or faults, thus representing the overhead.
$Rate_{failure}=\left(\frac{N_{failures}}{T}\right)\times 100$ (1)
$Rate_{fault}=\left(\frac{M_{faults}}{N_{failures}}\right)\times 100$ (2)
$Cost=100-Rate_{failure/fault}$ (3)
To statistically compare REST API testing tools for RQ1 and RQ3, we used non-parametric tests adhering to the recommended guidelines by Arcuri and Briand [60]. Specifically, we employed the Wilcoxon test to determine statistically significant differences, with a significance level set to 0.05 (i.e., $\alpha=0.05$). To further analyze which tool performed better, we calculated the effect size using the Vargha-Delaney Â12 measure. For RQ2 and RQ4, we present the results descriptively, focusing on discussions related to potential faults and regressions, similar to existing studies [27, 12].
Table 4: Number of test cases generated by different REST API testing tools for each API under test
RESTest-CBT
RESTest-ART
RESTest-RT
RESTest-FT
EvoMaster-BB
Schemathesis-PT
RESTler-F
RESTler-FL
RestTestGen-NET
RestTestGen-MAST
Alerts 400 400 400 401 7000 2101 731 392 800 0
Authentication 1308 1308 1308 1308 5511 2534 1892 1769 2440 0
Device-Karie 210 210 210 210 903 101 649 649 800 0
Device-Medido 0 455 0 0 9377 1837 860 860 1000 0
Device-Pilly 238 238 238 22 8802 102 395 398 400 0
Patients 545 545 545 471 5191 168 868 868 964 0
Measurements 510 510 510 510 5803 2136 4217 4734 1000 0
Users 574 0 574 0 8765 1363 476 476 3771 0
Courses 729 729 729 0 9740 78 423 418 0 0
Mobile App 528 528 528 528 9746 43 937 608 1020 0
Reimbursements 1182 1182 1182 1182 10393 370 1861 1861 1924 0
User Tasks 308 0 308 0 4719 733 255 255 2029 0
Invoicing 729 729 729 0 8448 493 1409 1405 1620 0
Catalog 592 592 592 592 11704 16 188 13 1060 0
Reports 689 689 689 301 8089 205 1129 1129 1512 0
Care 565 565 565 429 3160 6 20 2 600 0
Development 0 0 0 0 8592 9 1434 4 340 0
Total 9107 8680 9107 5954 125943 12295 17744 15841 21280 0
Table 5: RQ1 results: Unique failures categorized as 4XX and 5XX detected by each tool across all APIs from releases R1 to R14. Releases tagged with D, W, and M, denote daily, weekly, and monthly releases, respectively.
RT-CBT RT-ART RT-RT RT-FT EM-BB ST-PT RL-Fz RL-FL RTG-NET
Release 4XX 5XX 4XX 5XX 4XX 5XX 4XX 5XX 4XX 5XX 4XX 5XX 4XX 5XX 4XX 5XX 4XX 5XX
[W]R1 1198 4 1322 7 1209 6 1035 60 913 159 1229 31 169 23 92 27 142 31
[D]R2 394 7 531 4 381 5 569 61 4876 755 217 19 541 79 535 82 608 124
[D]R3 394 11 532 5 381 9 569 59 4876 751 217 13 541 78 535 73 610 119
[W]R4 394 9 532 6 381 5 569 61 4872 746 217 18 541 77 535 74 610 112
[D]R5 394 12 532 4 381 10 569 63 4875 752 217 996 541 79 535 82 610 126
[D]R6 394 10 532 7 380 10 569 61 4876 1633 217 11 541 80 535 79 610 118
[W]R7 394 10 532 6 381 6 569 59 4877 754 217 11 541 78 535 81 610 117
[D]R8 394 6 532 5 381 10 569 64 4878 758 217 15 541 79 535 80 610 117
[D]R9 394 9 532 8 380 10 568 64 4876 2810 217 12 540 83 535 80 610 122
[W]R10 394 9 532 5 381 6 568 61 4877 1156 217 12 541 81 534 82 610 122
[D]R11 394 12 532 7 381 8 568 65 4876 2090 217 22 541 77 534 78 610 128
[D]R12 395 6 533 4 382 3 662 61 5230 4059 218 12 541 86 535 83 797 115
[W]R13 395 3 533 2 382 7 663 58 5281 757 218 18 541 78 535 84 797 118
[M]R14 395 6 533 4 382 4 663 62 5282 1948 218 12 540 80 535 83 797 116
Table 6: Statistical comparison of various testing tools based on the number of failures detected. It includes the p-values derived from the Wilcoxon test and the $\hat{A}_{12}$ values obtained through Vargha-Delaney’s effect size measure.
RT-CBT RT-ART RT-RT RT-FT EM-BB ST-PT RL-Fz RL-FL RTG-NET
p-val $\hat{A}_{12}$ p-val $\hat{A}_{12}$ p-val $\hat{A}_{12}$ p-val $\hat{A}_{12}$ p-val $\hat{A}_{12}$ p-val $\hat{A}_{12}$ p-val $\hat{A}_{12}$ p-val $\hat{A}_{12}$ p-val $\hat{A}_{12}$
RT-CBT - - $<$0.05 0.557 0.779 0.5 $<$0.05 0.75 $<$0.05 0.036 $<$0.05 0.268 $<$0.05 0.268 $<$0.05 0.268 $<$0.05 0.268
RT-ART $<$0.05 0.443 - - $<$0.05 0.443 $<$0.05 0.74 $<$0.05 0.034 $<$0.05 0.268 $<$0.05 0.268 $<$0.05 0.268 $<$0.05 0.268
RT-RT 0.797 0.5 $<$0.05 0.557 - - $<$0.05 0.75 $<$0.05 0.036 $<$0.05 0.268 $<$0.05 0.268 $<$0.05 0.268 $<$0.05 0.268
RT-FT $<$0.05 0.25 $<$0.05 0.26 $<$0.05 0.25 - - $<$0.05 0.028 $<$0.05 0.256 $<$0.05 0.268 $<$0.05 0.268 $<$0.05 0.268
EM-BB $<$0.05 0.964 $<$0.05 0.966 $<$0.05 0.964 $<$0.05 0.972 - - $<$0.05 0.952 $<$0.05 0.733 $<$0.05 0.733 $<$0.05 0.76
ST-PT $<$0.05 0.732 $<$0.05 0.732 $<$0.05 0.732 $<$0.05 0.744 $<$0.05 0.048 - - $<$0.05 0.483 $<$0.05 0.483 $<$0.05 0.36
RL-Fz $<$0.05 0.732 $<$0.05 0.732 $<$0.05 0.732 $<$0.05 0.732 $<$0.05 0.267 $<$0.05 0.517 - - $<$0.05 0.716 0.186 0.499
RL-FL $<$0.05 0.732 $<$0.05 0.732 $<$0.05 0.732 $<$0.05 0.732 $<$0.05 0.267 $<$0.05 0.517 $<$0.05 0.284 - - 0.186 0.499
RTG-NET $<$0.05 0.732 $<$0.05 0.732 $<$0.05 0.732 $<$0.05 0.732 $<$0.05 0.24 $<$0.05 0.64 0.186 0.501 0.186 0.501 - -
* The values shown in gray color indicate that $p-value>\alpha$, i.e., there is no statistically significant difference. The values presented in bold indicate that the
tool listed on the left side outperformed the tool on the right side. Instances of self-comparison of tools are left blank in the table.
Figure 4: RQ1 results: Boxplot representing the median number of unique failures detected by each tool across all releases.
Figure 5: RQ2 results: Faults and their frequency of occurrence corresponding to each testing tool. Each fault is color-coded to highlight its uniqueness. The size of the bubble indicates the frequency of occurrence of each fault. Detailed information for these faults is presented in Table 8.
Figure 6: Regressions and their frequency of occurrence corresponding to each testing tool. Each type of regression is color-coded to highlight its uniqueness. The size of the bubble indicates the frequency of occurrence of each regression. Detailed information for all types of regressions is presented in Table 12.
Table 7: RQ2 results: Unique potential faults identified from the results generated by each tool across releases R1–R14. Releases tagged with D, W, and M, denote daily, weekly, and monthly releases, respectively.
RT-CBT RT-ART RT-RT RT-FT EM-BB ST-PT RL-Fz RL-FL RTG-NET
[W]R1 8 6 9 7 3 3 5 5 3
[D]R2 2 3 2 3 5 4 5 5 4
[D]R3 2 3 2 4 5 4 5 5 6
[W]R4 4 4 4 5 7 6 5 5 6
[D]R5 4 4 4 5 7 6 5 5 7
[D]R6 4 4 4 4 7 6 5 5 7
[W]R7 4 4 4 4 7 6 5 5 7
[D]R8 4 4 4 5 7 6 5 5 6
 Releases
[D]R9 4 4 4 5 6 6 5 5 6
[W]R10 4 4 4 5 6 6 5 5 7
[D]R11 4 4 4 5 7 6 5 5 7
[D]R12 3 3 2 4 6 5 5 5 6
[W]R13 2 2 3 3 5 5 5 5 6
[M]R14 3 3 2 4 5 5 5 5 6
Table 8: RQ2 results: Description of all potential faults, each associated with a specific fault ID and the HTTP status code, identified across all releases.
ID Status code Description
F1 502 & 413 Server failed to handle the request/response, potentially large data or invalid data/response from another server.
F2 200 Invalid login details lead to success with the empty response.
F3 500 Server failed to understand the request.
F4 400 Error in parsing request headers.
F5 429 Multiple invalid login attempts continuously increase login time for the next attempt.
F6 200 Incorrect user ID results in success.
F7 200 Empty/invalid JSON object for a device results in the success response with complete device information.
F8 409 Error while saving record based on ID which is not present in the database.
F9 200 Incorrect property values in JSON fetch all records for general tasks/surveys.
F10 500 Iterating over String type object results in server error.
F11 500 Database error when inserting records with missing fields.
F12 500 Server error when trying to schedule device’s update
F13 400 Type mismatch when performing arithmetic operation.
F14 500 Error when converting Python Int type large value to Long in C.
F15 400 Performing array operation in Boolean type variable.
F16 500 Accessing Dict items from String/None type object.
F17 409 Violation of foreign key constraint in the database.
F18 500 Failure in handling long value results in a complete database query as part of the response.
Table 9: RQ3 results: REST API coverage achieved by each testing tool for each release from R1 to R14. Releases tagged with D, W, and M, denote daily, weekly, and monthly releases, respectively.
RT-CBT RT-ART RT-RT RT-FT EM-BB ST-PT RL-Fz RL-FL RTG-NET
[W]R1 76.73% 78.7% 76.73% 77.62% 77.16% 77.16% 77.16% 77.16% 76.67%
[D]R2 81.18% 82.57% 81.21% 82.01% 83.81% 84.07% 84.04% 84.04% 83.93%
[D]R3 81.29% 82.67% 81.29% 82.06% 84.12% 84.13% 84.13% 84.12% 84.0%
[W]R4 81.3% 82.67% 81.3% 82.06% 84.14% 84.14% 84.14% 84.14% 84.01%
[D]R5 81.32% 82.67% 81.32% 82.06% 84.16% 84.16% 84.16% 84.16% 84.02%
[D]R6 81.32% 82.67% 81.32% 82.06% 84.16% 84.16% 84.16% 84.16% 84.02%
[W]R7 81.33% 82.67% 81.35% 82.06% 84.17% 84.18% 84.17% 84.17% 84.05%
[D]R8 81.35% 82.67% 81.35% 82.06% 84.18% 84.18% 84.18% 84.18% 84.05%
 Releases
[D]R9 81.35% 82.67% 81.35% 82.06% 84.18% 84.18% 84.18% 84.18% 84.05%
[W]R10 81.35% 82.67% 81.35% 82.06% 84.18% 84.18% 84.18% 84.18% 84.05%
[D]R11 81.35% 82.67% 81.35% 82.06% 84.18% 84.18% 84.18% 84.18% 84.05%
[D]R12 81.35% 82.67% 81.35% 82.06% 84.18% 84.18% 84.18% 84.18% 84.05%
[W]R13 81.36% 82.67% 81.36% 82.06% 84.19% 84.19% 84.19% 84.19% 84.07%
[M]R14 81.36% 82.67% 81.36% 82.06% 84.19% 84.19% 84.19% 84.19% 84.07%
Table 10: Statistical comparison of testing tools based on REST API coverage. It includes the p values calculated using the Wilcoxon test and the $\hat{A}_{12}$ values derived from the Vargha-Delaney effect size measure.
RT-CBT RT-ART RT-RT RT-FT EM-BB ST-PT RL-Fz RL-FL RTG-NET
p-val $\hat{A}_{12}$ p-val $\hat{A}_{12}$ p-val $\hat{A}_{12}$ p-val $\hat{A}_{12}$ p-val $\hat{A}_{12}$ p-val $\hat{A}_{12}$ p-val $\hat{A}_{12}$ p-val $\hat{A}_{12}$ p-val $\hat{A}_{12}$
RT-CBT - - 1.0 0.419 0.664 0.5 0.805 0.527 1.0 0.394 1.0 0.394 1.0 0.394 1.0 0.394 1.0 0.445
RT-ART $<$0.05 0.581 - - $<$0.05 0.581 $<$0.05 0.56 1.0 0.476 1.0 0.476 1.0 0.476 1.0 0.476 0.978 0.48
RT-RT 0.336 0.5 1.0 0.419 - - 0.805 0.527 1.0 0.394 1.0 0.394 1.0 0.394 1.0 0.394 1.0 0.445
RT-FT 0.195 0.473 0.997 0.44 0.195 0.473 - - 1.0 0.393 1.0 0.393 1.0 0.393 1.0 0.393 1.0 0.397
EM-BB $<$0.05 0.606 $<$0.05 0.524 $<$0.05 0.606 $<$0.05 0.607 - - 0.796 0.498 0.735 0.499 0.664 0.499 $<$0.05 0.527
ST-PT $<$0.05 0.606 $<$0.05 0.524 $<$0.05 0.606 $<$0.05 0.607 0.204 0.502 - - 0.336 0.501 0.265 0.501 $<$0.05 0.527
RL-Fz $<$0.05 0.606 $<$0.05 0.524 $<$0.05 0.606 $<$0.05 0.607 0.265 0.501 0.664 0.499 - - 0.415 0.5 $<$0.05 0.527
RL-FL $<$0.05 0.606 $<$0.05 0.524 $<$0.05 0.606 $<$0.05 0.607 0.336 0.501 0.735 0.499 0.585 0.5 - - $<$0.05 0.527
RTG-NET $<$0.05 0.555 $<$0.05 0.52 $<$0.05 0.555 $<$0.05 0.603 1.0 0.473 1.0 0.473 1.0 0.473 1.0 0.473 - -
* The values shown in gray color indicate that $p-value>\alpha$, i.e., there is no statistically significant difference. The values presented in bold indicate that the
tool listed on the left side outperformed the tool on the right side. Instances of self-comparison of tools are left blank in the table.
Table 11: RQ4 results: Number of potential regressions found with various testing tools across evolving releases, from R1 to R14.
RT-CBT RT-ART RT-RT RT-FT EM-BB ST-PT RL-Fz RL-FL RTG-NET
R1-R2 2 1 1 0 11 2 2 2 1
R2-R3 0 0 0 0 6 1 1 0 0
R3-R4 0 1 0 0 7 2 1 0 0
R4-R5 0 1 0 0 1 2 0 0 0
R5-R6 0 0 1 1 3 1 1 0 1
R6-R7 0 0 1 1 4 1 1 0 1
R7-R8 0 0 0 1 3 1 0 0 0
R8-R9 0 0 0 1 6 1 1 0 0
R9-R10 0 1 0 0 6 1 1 0 0
 Release changelog
R10-R11 0 1 0 0 2 2 1 1 0
R11-R12 1 0 0 0 4 2 1 2 0
R12-R13 1 0 0 0 3 2 1 1 0
R13-R14 0 0 0 0 2 2 1 1 0
Table 12: RQ4 results: Description of all potential regressions corresponding to each regression ID and release changelog.
ID Release changelog Description
Rg1 R1-R2, R9-R10 200 Ok to 500 Server error while iterating over String object
Rg2 R1-R2, R3-R4, R8-R9, R9-R10 200 Ok to 200 Ok with additional info or different responses
Rg3 R1-R2, R3-R4, R8-R9, R9-R10 200 Ok to 400 Validation error
Rg4 R1-R2 404 Not Found to 200 Ok
Rg5 R1-R2, R3-R4, R4-R5, R5-R6, R6-R7 502 Unsatisfied request to 400 Validation error
Rg6 R1-R2, R8-R9 400 Validation error to 200 Ok with additional details
Rg7 R1-R2, R2-R3 502 Unsatisfied request to 500 Server error
Rg8 R1-R2 502 Unsatisfied request to 200 Ok with additional details
Rg9 R1-R2, R6-R7, R12-R13 502 Unsatisfied request to 500 Server error caused by a database issue
Rg10 R1-R2, R2-R3 404 Not Authenticated to 200 Ok (with additional details/record)
Rg11 R1-R2 400 Error parsing headers to 400 Input Validation error
Rg12 R1-R2, R2-R3, R3-R4, R9-R10 400 Error parsing headers to 200 Ok with additional details
Rg13 R2-R3 200 Ok to 500 Server error caused by a database issue
Rg14 R2-R3 200 Ok to 404 Request URL not found
Rg15 R2-R3, R2-R3, R3-R4, R8-R9, R9-R10 500 Server error to 200 Ok with additional details
Rg16 R2-R3, R3-R4, R4-R5, R5-R6, R6-R7, R7-R8, R8-R9, R9-R10, R10-R11, R11-R12, R13-R14 400 Validation error (403 No access) to 502 Unsatisfied request
Rg17 R3-R4, R9-R10 500 Server error to 404 Request URL not found
Rg18 R3-R4 200 Ok to 500 Server error while accessing data from None type object
Rg19 R3-R4, R5-R6, R6-R7, R7-R8, R10-R11, R11-R12, R12-R13, R13-R14 500 Server error to 502 Unsatisfied request
Rg20 R3-R4, R10-R11 404 Not Found to 502 Unsatisfied request
Rg21 R4-R5, R6-R7, R7-R8, R8-R9, R11-R12, R12-R13, R13-R14 502 Unsatisfied request to 500 Server error
Rg22 R4-R5, R6-R7, R8-R9, R9-R10, R10-R11, R11-R12, R12-R13, R13-R14 502 Unsatisfied request to 404 Request URL not found (403 No access, 400 Input Validation error)
Rg23 R8-R9 404 Not Found to 500 Server error caused by a database issue
Table 13: RQ5 results: Failure and fault detection rates of tests generated by each tool, calculated as the percentage of failures detected and faults identified across all releases, from R1 to R14. Releases tagged with D, W, and M, denote daily, weekly, and monthly releases, respectively.
RT-CBT RT-ART RT-RT RT-FT EM-BB ST-PT RL-Fz RL-FL RTG-NET
Failure Fault Failure Fault Failure Fault Failure Fault Failure Fault Failure Fault Failure Fault Failure Fault Failure Fault
[W]R1 13.2% 0.67% 15.31% 0.45% 13.34% 0.74% 18.39% 0.64% 0.85% 0.28% 10.25% 0.24% 1.08% 2.6% 0.75% 4.2% 0.81% 1.73%
[D]R2 4.4% 0.5% 6.16% 0.56% 4.24% 0.52% 10.58% 0.48% 4.47% 0.09% 1.92% 1.69% 3.49% 0.81% 3.89% 0.81% 3.44% 0.55%
[D]R3 4.45% 0.49% 6.19% 0.56% 4.28% 0.51% 10.55% 0.64% 4.47% 0.09% 1.87% 1.74% 3.49% 0.81% 3.84% 0.82% 3.43% 0.82%
[W]R4 4.43% 0.99% 6.2% 0.74% 4.24% 1.04% 10.58% 0.79% 4.46% 0.12% 1.91% 2.55% 3.48% 0.81% 3.84% 0.82% 3.39% 0.83%
[D]R5 4.46% 0.99% 6.18% 0.75% 4.29% 1.02% 10.61% 0.79% 4.47% 0.12% 9.87% 0.49% 3.49% 0.81% 3.89% 0.81% 3.46% 0.95%
[D]R6 4.44% 0.99% 6.21% 0.74% 4.28% 1.03% 10.58% 0.63% 5.17% 0.11% 1.85% 2.63% 3.5% 0.81% 3.88% 0.81% 3.42% 0.96%
[W]R7 4.44% 0.99% 6.2% 0.74% 4.25% 1.03% 10.55% 0.64% 4.47% 0.12% 1.85% 2.63% 3.49% 0.81% 3.89% 0.81% 3.42% 0.96%
[D]R8 4.39% 1.0% 6.19% 0.74% 4.29% 1.02% 10.63% 0.79% 4.48% 0.12% 1.89% 2.59% 3.49% 0.81% 3.88% 0.81% 3.42% 0.83%
[D]R9 4.43% 0.99% 6.22% 0.74% 4.28% 1.03% 10.61% 0.79% 6.1% 0.08% 1.86% 2.62% 3.51% 0.8% 3.88% 0.81% 3.44% 0.82%
[W]R10 4.43% 0.99% 6.19% 0.74% 4.25% 1.03% 10.56% 0.79% 4.79% 0.1% 1.86% 2.62% 3.51% 0.8% 3.89% 0.81% 3.44% 0.96%
[D]R11 4.46% 0.99% 6.21% 0.74% 4.27% 1.03% 10.63% 0.79% 5.53% 0.1% 1.94% 2.51% 3.48% 0.81% 3.86% 0.82% 3.47% 0.95%
[D]R12 4.4% 0.75% 6.19% 0.56% 4.23% 0.52% 12.14% 0.55% 7.38% 0.06% 1.87% 2.17% 3.53% 0.8% 3.9% 0.81% 4.29% 0.66%
[W]R13 4.37% 0.5% 6.16% 0.37% 4.27% 0.77% 12.11% 0.42% 4.79% 0.08% 1.92% 2.12% 3.49% 0.81% 3.91% 0.81% 4.3% 0.66%
[M]R14 4.4% 0.75% 6.19% 0.56% 4.24% 0.52% 12.18% 0.55% 5.74% 0.07% 1.87% 2.17% 3.49% 0.81% 3.9% 0.81% 4.29% 0.66%
Table 14: RQ5 results: Testing cost overhead for each tool, calculated based on the percentage of failures detected and faults identified throughout all releases, from R1 to R14. Releases tagged with D, W, and M, denote daily, weekly, and monthly releases, respectively.
RT-CBT RT-ART RT-RT RT-FT EM-BB ST-PT RL-Fz RL-FL RTG-NET
Failure Fault Failure Fault Failure Fault Failure Fault Failure Fault Failure Fault Failure Fault Failure Fault Failure Fault
[W]R1 86.8% 99.33% 84.69% 99.55% 86.66% 99.26% 81.61% 99.36% 99.15% 99.72% 89.75% 99.76% 98.92% 97.4% 99.25% 95.8% 99.19% 98.27%
[D]R2 95.6% 99.5% 93.84% 99.44% 95.76% 99.48% 89.42% 99.52% 95.53% 99.91% 98.08% 98.31% 96.51% 99.19% 96.11% 99.19% 96.56% 99.45%
[D]R3 95.55% 99.51% 93.81% 99.44% 95.72% 99.49% 89.45% 99.36% 95.53% 99.91% 98.13% 98.26% 96.51% 99.19% 96.16% 99.18% 96.57% 99.18%
[W]R4 95.57% 99.01% 93.8% 99.26% 95.76% 98.96% 89.42% 99.21% 95.54% 99.88% 98.09% 97.45% 96.52% 99.19% 96.16% 99.18% 96.61% 99.17%
[D]R5 95.54% 99.01% 93.82% 99.25% 95.71% 98.98% 89.39% 99.21% 95.53% 99.88% 90.13% 99.51% 96.51% 99.19% 96.11% 99.19% 96.54% 99.05%
[D]R6 95.56% 99.01% 93.79% 99.26% 95.72% 98.97% 89.42% 99.37% 94.83% 99.89% 98.15% 97.37% 96.5% 99.19% 96.12% 99.19% 96.58% 99.04%
[W]R7 95.56% 99.01% 93.8% 99.26% 95.75% 98.97% 89.45% 99.36% 95.53% 99.88% 98.15% 97.37% 96.51% 99.19% 96.11% 99.19% 96.58% 99.04%
[D]R8 95.61% 99.0% 93.81% 99.26% 95.71% 98.98% 89.37% 99.21% 95.52% 99.88% 98.11% 97.41% 96.51% 99.19% 96.12% 99.19% 96.58% 99.17%
[D]R9 95.57% 99.01% 93.78% 99.26% 95.72% 98.97% 89.39% 99.21% 93.9% 99.92% 98.14% 97.38% 96.49% 99.2% 96.12% 99.19% 96.56% 99.18%
[W]R10 95.57% 99.01% 93.81% 99.26% 95.75% 98.97% 89.44% 99.21% 95.21% 99.9% 98.14% 97.38% 96.49% 99.2% 96.11% 99.19% 96.56% 99.04%
[D]R11 95.54% 99.01% 93.79% 99.26% 95.73% 98.97% 89.37% 99.21% 94.47% 99.9% 98.06% 97.49% 96.52% 99.19% 96.14% 99.18% 96.53% 99.05%
[D]R12 95.6% 99.25% 93.81% 99.44% 95.77% 99.48% 87.86% 99.45% 92.62% 99.94% 98.13% 97.83% 96.47% 99.2% 96.1% 99.19% 95.71% 99.34%
[W]R13 95.63% 99.5% 93.84% 99.63% 95.73% 99.23% 87.89% 99.58% 95.21% 99.92% 98.08% 97.88% 96.51% 99.19% 96.09% 99.19% 95.7% 99.34%
[M]R14 95.6% 99.25% 93.81% 99.44% 95.76% 99.48% 87.82% 99.45% 94.26% 99.93% 98.13% 97.83% 96.51% 99.19% 96.1% 99.19% 95.71% 99.34%
4.5 Results Discussion
Table 4 presents the number of tests each tool generated for every API. It can be observed that EvoMaster-BB consistently generated tests for every API and produced the greatest number of tests compared to all other tools. In addition, RestTestGen-NET, RESTler-F, RESTler-FL, and Schemathesis-PT were ranked second, and four variants of RESTest were ranked third in terms of the total number of tests generated. From the results, it is noticeable that RestTestGen-MAST did not generate any tests for all APIs even after successful runs. Given that RestTestGen-MAST is primarily designed to test for security vulnerabilities, it only generates tests upon identifying potential vulnerabilities (as mentioned in Github issue [^2] https://github.com/SeUniVr/RestTestGen/issues/17). This could potentially be attributed to the presence of read-only parameters in the API schema. However, we found that the APIs accessible to us do not contain such parameters. Alternatively, this could be related to the API implementation, which remains a black box to us due to our lack of access to the source code.
In subsequent subsections, we discuss results associated with failures, faults, coverage, regressions, and cost overhead corresponding to each RQ. For a discussion on the results, we refer to Tables 5 and 6 for RQ1 concerning failures, Tables 7 and 8 and Figure 5 for RQ2 pertaining to faults, Tables 9 and 10 for RQ3 associated with coverage, Tables 11 and 12 and Figure 6 for RQ4 related to regressions, and Tables 13 and 14 for RQ5 results concerning cost overhead. Note that while discussing the experimental results for each RQ, we maintain a balance between conveying essential information and omitting specific details to prevent disclosing sensitive information that could potentially affect the confidentiality and privacy of Oslo City’s industry partner.
4.5.1 RQ1 Results – REST API Failures
The API failures identified in our experiment results fall into two categories: client-side issues, represented by 4XX errors, and server-side issues, indicated by 5XX errors. From the results in Table 5, it can be observed that the test cases generated from all REST API testing tools resulted in a higher number of 4XX errors compared to 5XX errors across all 14 releases. In the case of RESTest, only the RESTest-FT generated tests led to a higher number of server-side failures compared to RESTest-CBT, RESTest-ART, and RESTest-RT. In particular, the test cases generated by EvoMaster-BB resulted in the highest number of 4XX errors and especially the maximum number of 5XX errors compared to all other testing tools. Since EvoMaster-BB generated the highest number of tests, detecting the highest number of failures could be attributed to this. Among Schemathesis-PT, RESTler-F, RESTler-FL, and RestTestGen-NET, the test cases generated by RestTestGen-NET lead to a higher number of client-side and server-side failures.
When comparing across releases R1 to R14, it is apparent that the tests generated by RESTest-CBT, RESTest-ART, RESTest-RT, RESTest-FT, and Schemathesis-PT resulted in a higher count of 4XX failures specifically for R1. However, for release R2, the number of 4XX failures was reduced. From R2 to R14, the tests derived from these tools produced a consistent number of client-side and server-side failures. For the tests generated by EvoMaster-BB, RESTler-F, RESTler-FL, and RestTestGen-NET, both 4XX and 5XX errors were relatively fewer for R1. However, from R2 to R14, the number of 4XX and 5XX errors was increased considerably. Furthermore, RESTler-F, RESTler-FL, and RestTestGen-NET demonstrated consistent results for both client-side and server-side failures across releases R2 to R14. On the other hand, EvoMaster-BB only exhibited consistent client-side failures from R2 to R14, while the number of server-side failures varied significantly across the same releases. Our observations revealed that a significant change in the number of failures from R1 to R2 was primarily due to a shift in the test environment. Release R1, in the staging environment, was transitioned from the prior release in the production environment. This environmental shift induced instability in R1, resulting in a higher detection of failures in the releases form from R2 onwards.
The aggregated results depicted in Figure 4 demonstrate that EvoMaster-BB detected the highest median number of unique failures compared to all other tools. Moreover, the median number of detected failures is comparable across all RESTest variants, with RESTest-ART and RESTest-FT marginally exceeding the others. For RESTler and RestTestGen, the median number of detected failures is nearly equivalent across their variants, i.e., RESTler-F, RESTler-FL, and RestTestGen-NET. In the case of Schemathesis-PT, the median number of detected failures is slightly higher than RESTler-F, RESTler-FL, and RestTestGen-NET, but lower than EvoMaster-BB and all variants of RESTest.
The results of the statistical comparison of all REST API testing tools, based on their ability to detect failures, are presented in Table 6. The comparative analysis reveals that RESTest-CBT outperformed both RESTest-ART and RESTest-FT. While RESTest-ART managed to outperform RESTest-FT, it falls short when compared to RESTest-RT. RESTest-FT, on the other hand, was unable to surpass any of the tools. Notably, EvoMaster-BB demonstrated superior performance over all REST API testing tools. In the case of Schemathesis-PT, it outperformed all variants of RESTest, including RESTest-CBT, RESTest-ART, RESTest-RT, and RESTest-FT. RESTler-F performed well compared to six tools, i.e., RESTest-CBT, RESTest-ART, RESTest-RT, RESTest-FT, Schemathesis-PT, and RESTler-FL. Lastly, RESTler-FL and RestTestGen-NET performed better than RESTest-CBT, RESTest-ART, RESTest-RT, RESTest-FT, and Schemathesis-PT.
[…]
4.5.2 RQ2 Results – REST APIs Faults
The results in Table 7 present the number of unique potential faults discovered from the results generated by each testing tool for all releases. It is noticeable that different numbers of potential faults were identified from the results of all tools for each release. Across all releases, the total number of faults found from the testing results of EvoMaster-BB and RestTestGen-NET was the highest compared to the other tools. Additionally, the results from Schemathesis-PT, RESTler-F, and RESTler-FL resulted in the second-highest number of potential faults. In the case of RESTest for R1, the most potential faults were identified from the results of RESTest-CBT and RESTest-RT.
Among all identified faults, some were common across tools, while others were unique to a specific tool due to the distinct testing techniques employed. Moreover, some faults were discovered from successful testing results (i.e., those with a 200 status code), while the rest were found due to client-side (4XX) and server-side (5XX) errors. In total, across all releases, 18 unique potential faults (F1-F18) were identified that are common across all tools, as detailed in Table 8. It was observed that certain faults reoccurred multiple times across evolving releases (i.e., from R1 to R14). The frequency of each fault’s occurrence w.r.t. each testing tool is illustrated in Figure 5.
The fault F1 is related to a server error, which was triggered when the server could not process requests containing invalid (e.g., missing required fields) or large amounts of data (e.g., string data exceeding 125KB). This fault was also noted in responses from third-party applications. Consequently, it is associated with two HTTP status codes: 502 Bad Gateway and 413 Payload Too Large. As depicted in the results (Figure 5), fault F1 was identified multiple times from the results of all testing tools. This fault was found in multiple APIs including Alerts, Authentication, Device-Karie, Device-Medido, Device-Pilly, Measurements, Users, Courses, Mobile App, Reimbursements, Invoicing, and Reports.
Fault F2 was detected from successful results that returned a 200 status code. This fault was observed in the Authentication API, where requests containing invalid login information resulted in 200 Ok responses. This implies that the system’s authentication process may not be functioning properly, permitting unauthorized access and thereby posing a considerable security risk. Therefore, F2 could potentially be classified as a security fault.
Fault F3 was also discovered in Authentication API, resulting in 500 Internal Server Error with a message stating that the server could not understand the request. This fault was discovered only through the results of RESTler-FL. It was noted that the API could not interpret the tokens generated by RESTler-FL for device authentication. Similarly, fault F4 was identified from errors related to 400 Bad Request, accompanied by a message stating that the server could not parse the headers. This fault was discovered from tests generated by RESTler-F and RESTler-FL. This fault was found in APIs including Alerts, Authentication, Device-Karie, Device-Medido, Measurements, Users, Invoicing, and Reports.
Fault F5 was discovered from the error 429 Too Many Requests originating from Authentication. This fault was detected in the tests generated by RESTest-ART, RESTest-RT, and RESTest-FT. In this case, it was observed that even after reaching the maximum limit for login attempts, the waiting time for the next attempts continued to increase from seconds to minutes to hours. This pattern suggests a potential fault in the API’s implementation, which is intended to control or limit the incremental wait time.
Fault F6 was found in Measurements API using the tests generated by RESTest-ART, RESTest-CBT, and RESTest-RT. This fault was identified when incorrect user IDs resulted in successful responses with 200 Ok status. Similarly, fault F7 was found in two device-related APIs, i.e., Medido and Pilly, and using the tests generated by the four variants of RESTest. In this case, empty/invalid JSON objects for these devices led to 200 Ok responses that contained complete information about the corresponding devices. Furthermore, a similar case was observed for the fault F9 that was identified in Users API using tests generated by RESTest-CBT and RESTest-RT. It was noticed that JSON objects with incorrect property values led to 200 Ok responses, providing all user-relevant records such as tasks/surveys. These faults (i.e., F6, F7, and F9) could be due to missing input validations in the API implementation.
The faults F8, F11, F17, and F18 are associated with errors that originated from the database. Fault F8 was discovered in Measurements API from the tests generated by RESTest-CBT and RESTest-RT. This fault, indicated by 409 Conflict status code, was identified when the API implementation tried to insert measurement records for users with IDs that do not exist in the database. The fault F11 was found in Users API from the status code 500 Internal Server Error and from tests generated by EvoMaster-BB, Schemathesis-PT, and RestTestGen-NET. This fault was identified from a database error when API implementation tried to insert records with missing fields. Similarly to F8, fault F17 was also found in Measurements API due to 409 Conflict error from tests generated by RESTest-CBT and RESTest-RT. While the status code for both faults is the same, the detailed response for F17 indicated a violation of foreign key constraints in the database. This violation occurred when trying to retrieve measurement records using invalid foreign keys. We have marked this fault as unique and have left it to the API development team to determine if they are the same. For fault F18, identified in the Mobile App API, a 500 Internal Server Error occurred due to long values in the request object. In addition, the error response was accompanied by a complete database query. This fault was detected using the tests generated by EvoMaster-BB, RESTler-FL, RESTler-F, RESTest-FT, and RestTestGen-NET.
The faults F10, F13, F14, F15, and F16 are associated with programming errors or problems related to business logic, categories which align with the predefined classification [61]. Fault F10 was discovered due to the 500 Internal Server Error response originating from APIs such as Device-Karie, Device-Medido, Device-Pilly and Users. This fault occurred when the API implementation attempted to iterate over an object of String type. Fault F13 was identified due to a 400 Bad Request error in Measurements API. This fault pertains to a type mismatch issue encountered during arithmetic operations. The fault F14 was found due to 500 Internal Server Error response from Reimbursements API. The primary cause of this fault was the conversion of a large value from Python’s Int type to C’s Long type. The fault F15 was located due to a 400 Bad Request error in the Authentication API. This fault occurred during the execution of an array operation on a Boolean-type variable. Fault F16 was discovered due to 500 Internal Server Error response from the Users API. This fault occurred while attempting to access Dict items from an object of the String type. Among these faults, F10 was the most frequently occurring, and tests from all tools led to its detection. In the case of F13, it was only identified from tests generated by RESTest-CBT. F14 was discovered from tests produced by both RESTler-FL and RESTler-F, while F15 was only located from the tests generated by RESTler-FL. In addition, F16 was detected using tests obtained from EvoMaster-BB, RestTestGen-NET, and Schemathesis-PT.
Finally, fault F12 was discovered in the Karie device API using tests generated by EvoMaster-BB, RESTest-ART, RESTest-CBT, RESTest-RT, RESTest-FT, Schemathesis-PT, and RestTestGen-NET. It was noted that a 500 Internal Server Error was triggered whenever a request was made to the device APIs while the device was in the middle of a scheduled update. While the device software was undergoing an update, this server error originated from the application’s API under test. Note that medical devices come from different vendors, and their software update schedules are unknown to us.
[…]
4.5.3 RQ3 Results – REST APIs Coverage
For release R1, the results (Table 9) show that all testing tools achieved an API coverage close to 77%. From R2 to R14, the API coverage for all tools increased to 84%. In the case of RESTest, it is noticeable that the API coverage attained by both RESTest-CBT and RESTest-RT is approximately 81%. Meanwhile, the coverage reached by RESTest-ART and RESTest-FT is slightly higher, around 82%, from the R2 release onwards. The remaining tools EvoMaster-BB, Schemathesis-PT, RESTler-F, RESTler-FL, and RestTestGen-NET achieved a nearly consistent coverage of approximately 84% across all releases from R2 through R14.
Upon a detailed analysis of the coverage results generated by the Restats tool, we observed that all REST API testing tools easily achieve nearly 100% path coverage, operation coverage, request and response content type coverage, parameter coverage, and parameter value coverage [^3] A detailed description of each type of coverage is provided by Corradini et al. [59].. However, these tools could not achieve 100% status code coverage and status code class coverage. As a result, their overall coverage reaches a maximum of only 84%. Status code coverage is related to comparing returned status codes with those defined in the API schema [59]. This implies that the tests generated by these REST API testing tools could not cover all the documented status codes. Similarly, status code class coverage refers to the ratio of success codes (2XX) and failure codes (4XX and 5XX) covered by tests [59]. This suggests that the tests generated by these REST API testing tools struggled to maintain a balance among different types of status code classes. It is also evident from the RQ1 results Section 4.5.1 that tests generated by all tools resulted in a higher number of 4XX errors than 5XX errors. The primary reason for the lower coverage of success codes (2XX) is that these tools often generated invalid inputs, resulting in more failures than successes. It is important to note that if a tool detects a higher number of failures, it may be due to the tool’s testing techniques being designed to uncover more failures.
The statistical comparison of the API test coverage, achieved by all the REST API testing tools, is presented in Table 6. It is noticeable that RESTest-CBT, RESTest-RT, and RESTest-FT were unable to outperform any of the testing tools. On the other hand, RESTest-ART outperformed RESTest-CBT, RESTest-RT, and RESTest-FT, indicating that, in terms of coverage, RESTest-ART perform better compared to the other variants of RESTest. Furthermore, EvoMaster-BB, Schemathesis-PT, RESTler-F, and RESTler-FL outperformed all four variants of RESTest and RestTestGen-NET, but exhibited comparable performance to each other. Lastly, in the case of RestTestGen-NET, it only demonstrated better coverage compared to RESTest-CBT, RESTest-ART, RESTest-RT, and RESTest-FT.
[…]
4.5.4 RQ4 Results – Regressions
The results for potential regressions identified between two subsequent releases, using the tests generated by all testing tools, are presented in Table 11. It is evident that the tests from EvoMaster-BB and Schemathesis-PT led to finding a higher number of regressions in all release changelogs. The tests generated by RESTler-F resulted in locating regressions in nearly all release changelogs, with the exceptions of R4-R5 and R7-R8. In the case of RESTest, the tests from RESTest-ART led to finding more regressions compared to RESTest-CBT, RESTest-RT, and RESTest-FT. Moreover, RESTler-FL only performed better than all variants of RESTest. Both RESTest-RT and RestTestGen-NET showed equivalent performance in identifying only three regressions in the same set of release changelogs. Overall, EvoMaster-BB’s test results led to the detection of the highest number of regressions compared to all other tools.
In total, 23 unique potential regressions (Rg1-Rg23) were identified using the tests generated by all tools. It was observed that certain regressions appeared in multiple release changelogs, while others were unique to particular tools. Table 12 outlines all identified regressions corresponding to different release changes. In addition, it was noted that certain regressions reappeared several times, from R1 to R14. The frequency of each regression’s occurrence, corresponding to the tools whose tests revealed these regressions, is depicted in Figure 6.
The 23 regressions can be classified into four categories:
  1. 1.
    Success-to-Failure: APIs that initially returned a successful 200 status code, but in the next release, resulted in a 4XX/5XX error response. This implies that a functionality working in one release has failed in the next. The regressions classified under this category include Rg1, Rg3, Rg13, Rg14, and Rg18.
  2. 2.
    Failure-to-Success: APIs that initially returned a 4XX/5XX error response but led to a successful 200 status code in the subsequent release. This suggests that functionality not working (potentially due to a bug) in one release has been fixed in the next release. The regressions Rg4, Rg6, Rg8, Rg10, Rg12, and Rg15 belong to this category.
  3. 3.
    Success-Enhancement/Reduction: APIs that returned a successful 200 status code in both the initial and subsequent releases, but the response in the next release exhibited new enhancements or reductions. The only regression that falls into this category is Rg2.
  4. 4.
    Failure-Shift: APIs that initially returned one type of client/server-side error but later returned a different type of error. This indicates a shift in the type of failures. The regressions Rg5, Rg7, Rg9, Rg11, Rg16, Rg17, Rg19, Rg20, Rg21, Rg22, and Rg23 belong to this category.
In the Success-to-Failure category, Rg1 was found in the User Tasks API, where a 500 server error occurred while iterating the String object for the subsequent release. This regression could potentially be related to the fault F10, as discussed in the RQ2 results (Section 4.5.2). Rg1 was observed in the release changelogs of R1-R2 and R9-R10 through the tests generated by RESTest-CBT, RESTest-RT, and EvoMaster-BB. The regression Rg3 was identified due to 400 validation errors in Assignments and Survey endpoints of User Tasks API. Rg3 was noticed in the changelogs of R1-R2, R3-R4, R8-R9, and R9-R10 using the tests generated by the RESTest-ART and EvoMaster-BB. The regression Rg13 was discovered in the User Tasks API, where a 500 server error was caused due to a database issue. The regression Rg14 was identified in User Tasks API due to 404 errors (URL not found), signifying a shift in the endpoints. The regression Rg18 was noted when 500 server errors occurred while accessing data from the None type object. This regression could potentially be associated with the F16 fault, as discussed in the RQ2 results (Section 4.5.2). Both Rg13 and Rg14 were identified in the R2-R3 release changelog, while Rg18 was observed in the R3-R4 release changelog. Rg13, Rg14, and Rg18 were found in tests generated by EvoMaster-BB.
In the category Failure-to-Success, several regressions were identified in various APIs. Rg4 was discovered in the endpoints of general tasks of the User Tasks API, where errors ‘URL not found’ were fixed, leading to 200 Ok responses. Rg6 was found in assignments and survey endpoints of User Tasks API, where validation errors were fixed resulting in 200 Ok responses. Rg8 was located in the Measurements API, where additional details were provided in the 200 Ok responses, indicating both a fix and an enhancement. Rg10 was identified in logs, assignments, and survey endpoints of Users and User Tasks APIs due to ”Not Authenticated” errors were resolved that led to 200 Ok responses with additional records. Rg12 was detected in assignments-related endpoints of User Tasks API in which parsing headers errors were fixed. Rg15 was identified in logs and general tasks endpoints of User and User Tasks APIs, in which 500 server errors were fixed and resulted in 200 Ok responses containing refined and additional information. Furthermore, the regressions Rg4 and Rg8 were observed in the R1-R2 release changelogs. Rg6 appeared in the changelogs for the releases R1-R2 and R8-R9, while Rg10 was noted in the releases R1-R2 and R2-R3. Rg12 was found in the release changelogs of R1-R2, R2-R3, R3-R4 and R9-R10. Rg15 was located in the release changelogs of R2-R3, R3-R4, R8-R9, and R9-R10. The regressions Rg4, Rg6, Rg8, Rg12, and Rg15 were identified from tests generated by EvoMaster-BB. Furthermore, Rg8 was discovered through tests produced by both EvoMaster-BB and RESTler-FL.
In the Success-Enhancement/Reduction category, Rg2 was observed in the Users API related to the message logs endpoint. This regression was identified in the changelogs for releases R1-R2, R3-R4, R8-R9, and R9-R10, demonstrating multiple changes in users’ message logs during evolution. Moreover, this regression was found through the tests generated by RESTest-CBT and EvoMaster-BB.
Many regressions, classified as Failure-Shift, signify unstable changes that occur during API evolution and lead to different failures. Among these regressions, Rg17, Rg19, Rg20, and Rg23 are related to Development API, indicating continuous modifications in the API. The regressions Rg5, Rg11, Rg16, Rg20, and Rg22 are associated with search methods (e.g., search with name or ID) of Users API. In addition, Rg16 and Rg22 were identified in the Reports API during the generation or retrieval of report data. The regression Rg9 was discovered in the subscriptions-related endpoints of the Invoicing API. Regressions Rg7 and Rg21 were discovered in APIs related to Device and Users, indicating server errors when iterating over String-type objects. This regression could potentially be related to the F10 fault (RQ2 results Section 4.5.2). Furthermore, it can be observed (from Figure 6) that the regressions Rg16, Rg19, Rg21, and R22 were the most frequently occurring across multiple release changelogs.
Interestingly, regressions in this category were detected from tests obtained from nearly all testing tools. For instance, regression Rg5 was detected through tests generated by RESTest-ART, EvoMaster-BB, Schemathesis-PT, and RESTler-F. Similarly, regression Rg7 was identified by tests generated by tools EvoMaster-BB, Schemathesis-PT, RESTler-F, and RESTler-FL. On the other hand, regressions Rg9, Rg17, and Rg23 were only found using tests obtained from EvoMaster-BB, while Rg11 was specifically identified from the tests produced by RESTler-F. The regression Rg16 was detected by a variety of testing tools including RESTest-ART, RESTest-FT, EvoMaster-BB, Schemathesis-PT, RESTler-F, and RESTler-FL. Both Rg19 and Rg21 were identified by tests from tools RESTest-CBT, RESTest-RT, EvoMaster-BB, Schemathesis-PT, RESTler-FL, and RestTestGen-NET. The regression Rg20 was found using tests obtained from EvoMaster-BB and Schemathesis-PT. Lastly, regression Rg22 was located using tests generated by RESTest-ART, RESTest-FT, EvoMaster-BB, Schemathesis-PT, RESTler-F, and RESTler-FL.
[…]
4.5.5 RQ5 Results – Cost Overhead
Table 13 presents the results for failure and fault detection rates of tests generated by each tool throughout all releases. Additionally, Table 13 presents the results for the testing cost overhead associated with each tool across all releases. For RESTest-CBT and RESTest-RT, the failure rate is around 13% for R1 and 4% for R2 to R14, resulting in a cost overhead of approximately 86% for R1 and 95% for R2 to R14. Moreover, the fault detection rate for both RESTest-CBT and RESTest-RT varies between 0.5 and 1.0% for all releases, resulting in a roughly 99% cost overhead. In the case of RESTest-ART, the failure rate is slightly high around 15% for R1 and approximately 6% for other releases, resulting in a cost overhead of 84% for R1 and 93% for others. Similarly, RESTest-FT has a higher failure rate than other RESTest variants, about 18% for R1 and approximately 10-12% for other releases, leading to a cost overhead of 81% for R1 and 87-89% for others.
In the case of EvoMaster-BB, the failure rate is 0.85%, and the fault detection rate is 0.28% for R1. However, from R2 to R14, the failure rate increases to between 4% and 7%, while the fault detection rate decreases to approximately 0.1%. Consequently, the resulting cost overhead for failure cases also decreases from R2 to R14, approximating between 92% and 95%. However, the cost overhead for fault detection remains high, around 99%. Note that EvoMaster-BB generated the highest number of tests compared to other tools, hence its resulting ratios are lower. Similar is the case with Schemathesis-PT results. For R1 and R5, the failure rate is around 10% and the fault detection rate is up to 0.5%, resulting in a lower cost overhead of approximately 90% for only failure cases. For the remaining releases, the failure and fault rates are reduced to around 2%, thereby leading to a cost overhead of approximately 98%.
Interestingly, the results for RESTler-F, RESTler-FL, and RestTestGen-NET exhibit similar patterns. For R2 to R14, the failure and fault detection rates are approximately 3.5% and 0.8%, leading to cost overheads of 96% and 99%, respectively. For R1, the fault detection rates for RESTler-F, RESTler-FL, and RestTestGen-NET are 2.6%, 4.2%, and 1.7%, respectively, while the failure rates for these tools remain around 1%. Given its higher fault detection rate, RESTler-FL exhibits a lower cost overhead compared to the other tools.
In general, for all tools, the failure rate is higher in R1 compared to R2–R14. Consequently, the cost overhead is lower in R1 than in the remaining releases. However, only RESTler-F, RESTler-FL, and RestTestGen-NET have a higher fault detection rate in R1. For other tools, the fault detection rate increased from R2 to R14, as noticed with RESTest-CBT in R8 and RESTest-RT in R4–R11. These findings suggest that despite the high-cost overhead, the fault detection rate also increases in some cases. Furthermore, in the case of EvoMaster-BB, despite demonstrating a high-cost overhead, the results from Section 4.5.4 indicate that tests generated by EvoMaster-BB lead to the detection of a higher number of potential regressions. These results from cost overhead analysis indicate a trade-off among tools depending on the testing objective, i.e., failures, faults, coverage, or regressions. Hence, practitioners should first clearly define the testing objective, and then select the appropriate tool accordingly.
[…]
4.6 Threats to Validity
As with any experiments, potential threats to validity may arise in our experiments. In the following, we outline each potential threat that could impact our experimental results and the corresponding mitigation strategies.
4.6.1 External validity threat
The threat to external validity is related to the generalizability of experimental findings. A potential threat to the external validity of our experiments could arise from the REST APIs and tools selected for testing. In our experiments, we used 17 REST APIs with 120 endpoints from a real-world healthcare IoT application, along with its 14 evolving releases during DevOps. This experimental apparatus was provided by Oslo City’s healthcare department. It is worth noting that this application is operational across multiple municipalities in Norway, including Oslo, and is also being used in several other countries. Given the large scope of our application under test, the APIs utilized in our experiments constitute a representative sample. Moreover, we carefully selected REST API testing tools that align with the dynamics of our APIs under test. These tools, being state-of-the-art and well-established, make a good representative set. While our findings may not generalize to all healthcare applications, this limitation is common in empirical studies [13, 62].
4.6.2 Internal validity threat
The threat to interval validity can occur due to uncontrolled factors. A possible threat to the internal validity of our experiments may occur due to any instrumentation performed on the REST API testing tools or the REST APIs under test. To mitigate this threat, we selected REST API testing tools in their executable form, without modifying their code. In the case of REST APIs under test, we only have black-box access to these APIs, with all implementations carried out by the industry partner. Another possibility of internal validity threat can be due to tool inputs, including the API schema and test configuration, particularly when using RESTest. To reduce this threat, we relied on the documentation provided for each REST API testing tool to ensure proper setup and usage. Additionally, we conducted multiple sessions with the technical team from Oslo City and their industry partner. The primary purpose of these sessions was to demonstrate the setup process and acquire their feedback.
4.6.3 Construct validity threat
Construct validity threat is associated with the interpretation of results. We presented our evaluation findings using metrics commonly employed in similar experimental contexts [27, 9, 48]. Another potential threat to construct validity could arise from the manual analysis of faults and regressions. To mitigate this threat, we shared our experimental data, which includes generated test cases, failure reports, test coverage, regressions, and faults catalog, with Oslo City’s healthcare department for their review and feedback. Due to the sensitive nature of the data involved in this study, relating to API information and patient-associated data, we are bound by non-disclosure agreements that prevent us from publicly releasing a replication package or data set.
4.6.4 Conclusion validity threat
The threat to conclusion validity arises from the selection of analysis metrics, measures, or statistical tests that may influence the accurate drawing of conclusions. A potential threat to the validity of the conclusions of experiments may arise due to the randomness factor of REST API testing tools. We ran each tool 10 times with a fixed one-hour duration to address the randomness factor. This experimental setting is aligned with existing empirical studies [9, 48]. To analyze and report the results of our experiments, we employed metrics such as API failures, faults, test coverage, and regressions in evolving releases, which are commonly used in similar studies [27, 12]. Additionally, we evaluated the results using non-parametric statistical tests, specifically the Wilcoxon signed-rank test and Vargha-Delaney Â12 effect size measure, following recommended guidelines [60].
Figure 7: A workflow illustrating the recommendations for practitioners in selecting REST API testing tools.
5 Tool Selection Recommendations
Based on our experimental findings, we provide recommendations for practitioners to assist in selecting REST API testing tools, taking into account the testing goal and cost overhead. Figure 7 presents a workflow serving as a visual guide to illustrate the selection process. As observed in experimental results, there is no one-size-fits-all tool that is universally effective for all testing purposes. The performance and effectiveness of each tool can vary depending on the specific testing context. Therefore, testers need to define the testing goal, i.e., whether they are interested in analyzing failures, faults, coverage, regressions, or a combination of these aspects. In the following, we discuss recommendations considering the testing goal and cost overhead.
  • If the testing goal is to find the maximum number of unique failures and a high-cost overhead is acceptable, EvoMaster-BB can be a suitable choice. However, if the goal is to identify as many unique failures as possible with minimum cost overhead, RESTest-FT seems to be a reasonable alternative.
  • For detecting maximum unique potential faults, variants of RESTest such as RESTest-CBT, RESTest-RT, and RESTest-FT can be employed interchangeably. Nevertheless, if the priority is to keep the cost overhead as low as possible, RESTest-FT is an appropriate choice.
  • If the testing goal is to achieve maximum coverage and high-cost overhead is bearable, one can choose from a range of tools such as EvoMaster-BB, Schemathesis-PT, RESTler-F, RESTler-FL, and RestTestGen-NET. However, for reasonable coverage with minimum cost overhead, either RESTest-ART or RESTest-FT would be suitable.
  • When the testing goal is to discover a higher number of regressions across evolving releases and a high-cost overhead is tolerable, EvoMaster-BB can be an appropriate selection. On the other hand, if the requirement is for a low-cost overhead, Schemathesis-PT can be considered as an alternative.
  • If the testing goal is to identify the maximum number of unique failures and faults, and high-cost overhead is not a concern, one can opt for RESTest-CBT or RESTest-RT. However, if the requirement is to minimize cost overhead while detecting a higher number of unique failures and faults, RESTest-FT can be an appropriate choice.
  • When the testing goal is to identify the maximum number of unique failures while achieving high coverage, and if high-cost overhead is acceptable, EvoMaster-BB can be a suitable option. Conversely, if there is a need to reduce cost overhead, RestTestGen-NET can be viewed as a viable alternative.
  • If the testing goal is to detect the highest number of unique faults while ensuring high coverage with a sizable high-cost overhead, RestTestGen-NET could be an appropriate choice. However, if the same testing goal needs to be achieved with lower cost overhead, RESTler-F appears to be a suitable alternative.
  • When testing aims to achieve high coverage and detect the highest number of regressions across multiple releases, EvoMaster-BB can be utilized if a high-cost overhead is acceptable. On the other hand, if there is a need to reduce cost overhead while achieving the same testing goal, Schemathesis-PT appears to be a reasonable alternative.
  • If testing aims to uncover a maximum number of unique failures and regressions without high-cost overhead constraints, EvoMaster-BB can be an appropriate selection. Conversely, if the intent is to reduce cost overhead while maintaining the same testing goal, Schemathesis-PT can be considered a suitable option.
  • If the testing goal is to identify the maximum number of unique faults and regressions, and if a high-cost overhead is tolerable, EvoMaster-BB could be an appropriate choice. However, if the goal is to maintain low costs while still uncovering maximum unique faults and regressions, RESTler-F can be selected as an alternative option.
6 Experiences, Insights, and Challenges
Based on our experiments with REST API testing tools, we discuss experiences, insights, and challenges in the following sections.
6.1 Language-specific Test Inputs
The healthcare application used in our experiments operates in several countries and provides multilingual support. This trait is common among web services of multinational companies, which typically support multiple languages. An important aspect of testing REST APIs of such applications is using test inputs in all supported languages, necessitating the generation of language-specific test inputs. For instance, if an application supports Norwegian, Swedish, and English, generating test inputs in these languages is needed to test the application’s APIs from the language perspective. Moreover, if an application is region-specific with API supporting a particular language, it would require testing tools to generate language-specific inputs. For example, the cancer registry of Norway receives and compiles cancer cases in the Norwegian language [63]. In this case, the testing tool must generate test inputs in Norwegian.
The REST API testing tools used in our experiments mainly generate test inputs in English. Therefore, augmenting these tools to support the generation of test inputs specific to various languages emerges as a significant factor. Another potential avenue to explore is the automatic recognition of supported languages by these tools, followed by the generation of test inputs tailored to each identified language. Given the wide range of applications and rising popularity of large language models (LLMs), generating multilingual test inputs with LLMs could be a valuable enhancement to REST API testing tools. This approach would broaden the applicability of these tools in a multilingual context.
6.2 API Keys and Access Tokens
Healthcare IoT applications, such as the one in our study, are integrated with different third-party applications using API keys and access tokens for authentication and authorization. Each of these third-party applications supports different types of tokens (for example, OAuth2) with a certain level of security. Generating test cases for APIs that require access to third-party applications involves generating API keys and access tokens, which must adhere to a specific format. The automatic generation of valid API keys and access tokens presents a significant challenge.
6.3 Automation Level
Each REST API testing tool supports different automation levels. In addition to adhering to the OpenAPI specification, some tools require additional manual effort. For instance, RESTest requires writing test configurations. This effort escalates considerably as the REST APIs and their specifications evolve during the DevOps process. A potential solution to alleviate manual effort involves utilizing automated tools that can infer REST API specifications as their implementation undergoes evolution. For example, a tool like RESTSpecIT [64] can be used for this purpose.
6.4 Regression Test Selection, Minimization, and Prioritization
Numerous techniques are presented in the literature for prioritization, minimization, and selection of the test suite in the context of regression testing [53, 54]. However, to the best of our knowledge, their applicability to prioritize, minimize, and select regression test cases for REST APIs remains undetermined. Proposing new techniques for regression test suite prioritization, minimization, or selection falls outside the scope of this paper. Therefore, in our experiments, we used retest-all strategy, which involves running all the tests generated for the initial release on all subsequent releases. Given the daily or weekly release routine within the DevOps process, executing all test cases becomes impractical due to the significant overhead of API calls. For IoT applications’ REST APIs dependent on various third-party applications, the frequent execution of numerous tests is infeasible due to the API call limit imposed by these external applications. A potential area for future research could focus on devising strategies for prioritization, minimization, or selection of test cases specifically for the regression testing of REST APIs of IoT-based applications. The regression testing outcomes presented in this work could serve as a foundation for further research. These findings could also be used to refine or customize existing methods for test selection, minimization, and prioritization within the REST API testing context.
6.5 REST API Regression Testing Tool
Regression testing of continuously evolving REST APIs necessitates a specialized tool that supports automation. Until now, approaches and tools targeting REST API testing have primarily focused on OpenAPI specifications, REST API test generation and execution, or regression testing. Using separate tools for each task in the regression testing of evolving REST APIs can lead to significant manual effort and compatibility challenges. Therefore, one potential direction worth exploring is the creation of a toolchain that leverages existing tools designed for individual issues. This approach could streamline the process and enhance efficiency in the regression testing of evolving REST APIs.
7 Conclusion
Given the numerous available REST API testing tools, a key concern of the Oslo City Healthcare Department is selecting a suitable tool to test real-world IoT applications in healthcare in a DevOps environment. Taking this into account, we conducted an empirical study involving five well-established and state-of-the-art REST API testing tools: RESTest, EvoMaster, Schemathesis, RESTler, and RestTestGen. These tools were employed to test an industrial healthcare IoT application comprising 17 REST APIs with 120 endpoints. The experiments were carried out using 14 releases (including daily, weekly, and monthly) that experienced continuous development and evolution during the DevOps process. Given the black-box access to the APIs of our application under test, we set up black-box testing techniques supported by REST API testing tools. These configurations resulted in a total of 10 tools, namely RESTest-CBT, RESTest-ART, RESTest-FT, RESTest-RT, EvoMaster-BB, Schemathesis-PT, RESTler-F, RESTler-FL, RestTestGen-NET, and RestTestGen-MAST. Our experiments focused on analyzing failures, faults, test coverage, regressions in REST APIs under test, and cost overhead.
The results revealed that EvoMaster-BB generated the highest number of test cases compared to all other tools. On the other hand, RestTestGen-MAST did not generate any tests, indicating that none of the APIs has security vulnerabilities. In terms of API failures, all tools generated a higher number of client-side errors than server-side errors. Among all, EvoMaster-BB generated the highest number of errors on the client and server side. The tests produced by all tools led to the identification of 18 unique potential faults across 14 releases, with RESTest-CBT and RESTest-RT being the primary contributors to these faults. As for API test coverage, the results indicated that all tools attained a coverage level of 82-84%, with the lowest coverage observed in status code and status code class coverages. The regression results showed that tests generated by EvoMaster-BB led to finding the highest number of regressions. In total, 23 unique potential regressions were identified in all 14 releases. Finally, the cost analysis demonstrated that each tool generates a significant number of tests, often leading to numerous duplicate failures, thereby resulting in a cost overhead exceeding 80%.
Acknowledgements
This research work is a part of the WTT4Oslo project (No. 309175) funded by the Research Council of Norway. All the experiments reported in this paper are conducted in a laboratory setting of Simula Research Laboratory; therefore, they do not by any means reflect the quality of services Oslo City provides to its citizens. Moreover, these experiments do not reflect the quality of services various vendors provide to Oslo City.